OpenStudy (anonymous):

Help Me!! The volume (V) of a box is V = lwh. Find w, if V = 0.504, l = 0.9, and h = 0.7. Answers A) w = 8 B) w = 0.8 C) w = 0.08 D) w = 80

OpenStudy (kamille):

\[V=lwh\] \[w=\frac{ V }{ lh }\] all you need to do now is insert numbers!

OpenStudy (anonymous):

Just simple division here: V = lwh 0.504 = 0.9*w*0.7 = 0.63w w = 0.504/0.63 = 0.8 w = 0.8 ---> answer B